The inner diameter of a cylindrical wooden pipe is 24 cm
an.a. its outer diameter is 28 cm. The kngth of the pipe is
35 cm. Find the mass of the pipe. if 1 ems of wood has a
mass of 0.6 g.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Volume of the wood = Jt(r 2 - ri>h = 2: x ((14>2 - (12>2J x 35 ems SURFACE AREAS AND VOLUMES 197 = 110(196 - 144) cm3 = 5720 cm3 Mass of the pipe (or wood) = 0.6 >< 5720 g = 3432 g = 3.432 kg.